AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
ISOLDE
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Nov 15 21:31:58 2025
Author
*
:
Subject
*
:
> Transcription of paper logbook for runs. Please check as I cannot read all the writing. > > Wednesday 13/9/2017 > > Detectors installed: > 2083-9, 74 um, dE layer > 3080-23, 1012 um, E layer > 2410-23, 65 um, HI layer > > Det. Bias Leak > dE -24 0.46 > E -350 2.61 > HI -14 1.05 > > Runs 0 - 5 : junk > > Run 10 > Header: Alpha calibration > Date: 13/9/2017 > Start: 18:03 (CERN) > Stop: 18:12 > Note: Some cables disconnected to test CANNOT READ > > Run 11 > Header: Alpha calibration > Date: 13/9/2017 > Start: 18:16 > Stop: > Notes: > ADC Ch Notes > 1 0-15 ok > 1 16-31 > 2 0-15 7+8 only pulser > 2 16-31 ok > 3 0-15 only pulser > 3 16-31 19 low gain (other notes but crossed out) > 4 0-15 6 low (noise?), rest fine > 4 16-31 only pulser > 5 0-15 "probably okay" > 5 16-31 "probably okay" > 6 0-31 "probably okay" > > ------- > Thursday 14/9/2017 > > Det. Bias Leak > dE -24 0.44 > E -351 2.4 > HI -14 0.95 > > Run 12 > Header: > Date: 14/9/2017 > Start: 11:50 > Stop: 12:08 > Note: Bad. Pulser switched on through run > > Run 13 > Header: Pulser ON > Date: 14/9/2017 > Start: 12:07 > Stop: > Notes: good > ADC Notes > 1 ch 0 ... bad resolution in pulser > 2 ch 7,8 no alphas > 3 ch 0 low gain, ch 0-15 almost only pulser but maybe shieled from alpha source & low noise > 3 ch 24 low resolution > 4 ch 6 no alphas, ch 7 low gain for alphas but not for pulser > 4 ch 16-31 no alpha (maybe shielded & low noise) > 5 okay > 6 okay > > TDC missing #1-32, #1-80,120,123 > > 14:00 before turning off biases: > > Det. Bias Leak > dE -24 0.4 > E -351 2.15 > HI -14.5 0.89 > > 15:15 - Vent to remove dE layer > > Extra SSD cable arrives > > TDC mapping > > Det. Side TDC ch > dE p+n 16-63 > E n+n 64-79 > HI p+n 80-127 > > Bias detectors > Det. Bias Leak > E -350 1.83 > HI -14 0.8 > > Run 15 > Header: > Date: 14/9/2017 > Start: 18:22 > Stop: 18:25 > Notes: Bad. Pulser to 2 Hz > > Run 16 > Header: Alpha source in front of E detector > Date: 14/9/2017 > Start: 18:31 > Stop: > Notes: Bad. CANNOT READ > > Run 17 > Header: > Date: 14/9/2017 > Start: 18:32 > Stop: > Notes: Bad. ADC #4 ch 6, 7 bad. > Switch U7e8 (??) (n+n side p+n) from E detector > Problem moved -> CANNOT READ > > ------- > Friday 15/9/2017 > > 10:00 bias detectors > > Det. Bias Leak > E -351 2.13 > HI -14.5 0.88 > > T: -4.9 C > P: 4.9 x 10^-6 mbar > > Run 19 > Header: alpha-calib of E layer 3080-23 > Date: 15/9/2017 > Start: > Stop: > Notes: > No TDC data ...? > ADC Ch Comments > 1 0-31 Pulser (only) in all channels except ch2 which has a few counts across whole spectrum (hair) > Resolutions range from 2.0 -> 7.0 > Pulser centroid ~ channel 2800 in all spectra (except ch 2 of course) > 2 0-15 Pulser in all channels, ~channel 2000 (i.e. similar gains) > Resolution ~ 2.0 ch, except ch 8 sigma = 4.5 ch > 2 16-31 Pulser in all channels at ~channel 3000 > Resolution ~ 2.0 ch, excehpt ch 24 which is 6.7 ch & ch 16: 4.35 ch. > 3 0-31 Pulser & alphas in all channels > (Not 19...see below) > Resolution is *poor* in ch0 > All gains look similar > Alpha spectra in similar channel numbers as for other ADCs, but pulser at only 1/2 the channel number > (A drawing is given to compare ch0 to ch1-15.) > 3 19 Gain seems very low > Pulser @ channel 430, alphas @ channel 200 > So about 1/5 of gain of others in ADC 3 channels > 4 0-15 All but ch 6 very similar (good) > ch 6: Pulser in similar channel to other spectra, but narrower (~4 ch, c.f. ~8-10 channels for other spectra.) > And alphas are much lower gain, appearing ~ ch 70, c.f. 1100 for others > 4 16-31 All okay. Triple alpha @ ~ channel 1150, pulser @ 2800 > 5 0-31 All okay (though low gain - deliberately) > 6 0-31 Ok, see above for ADC 5 notes. > > Det. Bias Leak > E -351 2.08 > HI -14.5 0.88 > > Restart MIDAS; VME crate > => TDC signals return > > 12:02 Removed E layer #3080-23 (HI layer #2410-23 remains unchanged) > > New 1035 um Si installed #2951-22. It uses a different bias voltage than the previous E layer (-160 vs. -351 V) > > 12:50 Stop DAQ (run 22 larger than necessary) > > Run 23 (?? no note) > Header: > Date: > Start: > Stop: > Notes: > ADC ch Comments > 1 0-15 Pulser only, similar gains > ch 2 no pulser (hair) > No alphas (unsurprising!) > 1 16-31 Pulser only, similar gains > All okay. > (No alphas, unsurprising!) > 2 0-31 Pulser only, no alphas, similar gains > All okay > 3 0-15 Pulser + alphas > alphas in all strips > alpha similar gains and resolutions > pulser in ch 0 at ~1/2 amplitude > 3 16-31 Pulser + alphas. > ch 19 pulser and alphas at much reduced channel number ~1/5 gain > 4 0-15 Pulser + alphas > ch 6 no alphas - Pulser narrow (no detector capacitence?) > ch 7 alphas at reduced gain (~1/5), pulser narrow > 4 16-31 All good > 5 0-31 All good (low gains - by design) > 6 0-31 All good (low gains - by design) > > v767 #1 TDC > ch Comments > 0 1 peak > 1-15 No data > 16-31 1 peak > 32 No peaoks > 33 1 peak + 1 small peak > 34-44 2 peaks > 45-47 1 peak > 48 1 peak + hair > 49-52 1 peak > 53 1 peak + hair > 54-56 2 peaks > 57-63 1 peak > 64-79 broad region of data (alphas?) > 80-86 1 peak > 87 1 peak + 1 smaller peak > 88 Mainly 1 or 2 peaks + lots of hair > 89 1 peak > 90,91 2 peaks > 92-95 1 peak > 96 1 peak > 97-127 Mainly 1 peak > > Cable for 1 mm E detector must be faulty > > Before opening the chamber: > Det. Bias Leak > E -160 1.12 > HI -14.5 0.92 > > (When trying to switch the cable to the SSD, apparently it was not fully connected. Some time is lost to pump, vent, check, etc) > > After new cabling is confirmed and evacuation is complete: > Det. Bias Leak > E -160 1.17 > HI 14.5 0.96 > > Run 24 > Header: Check new cable > Date: 15/9/2017 > Start: 16:18 (UK) > Stop: 16:35 (UK) > Notes: > ADC ch Comments > 3 0-15 okay > 3 16-31 ch 19 low gain > 4 0-15 ch 5, ch 7 low gain > 4 16-31 okay
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.3-7933898